If P(x) is a fifth degree polynomial and Q(x) is a fourth degree polynomial, what can be said about the degree of P(x) - Q(x)? What I need is an example of this. Thank you
Answer by ilana(307)

Here's an example:
P(x)=x^5+x^4+x+2
Q(x)=x^4+1
P(x)-Q(x)=x^5+1
Can you tell from this what the degree of P(x)-Q(x) will always be?